OpenStudy (xlpoisonpotato):

Two boats start their journey from the same point A and travel along directions AC and AD, as shown below: ABC is a right triangle with measure of angle ABC equal to 90 degrees and length of AB equal to 400 feet. There is a point C on BD such that measure of angle ACB is 60 degrees and measure of angle ADC is 30 degrees. What is the distance, CD, between the boats? 461.9 ft 530.9 ft 646.4 ft 325.5 ft

OpenStudy (xlpoisonpotato):

hey can i hget some help

OpenStudy (mecharv):

Where is the diagram?

OpenStudy (xlpoisonpotato):

hello?

OpenStudy (mecharv):

there is some problem with that link. please check

OpenStudy (xlpoisonpotato):

oh

OpenStudy (xlpoisonpotato):

it shows that its 400 feet tall

OpenStudy (xlpoisonpotato):

at one end away from the right angle it it 30 degrees

OpenStudy (xlpoisonpotato):

then closer to the 90 degree angle it has an angle of 60

OpenStudy (shadowlegendx):

Save the image, download it, then upload it here, using the "Attach File" function

OpenStudy (xlpoisonpotato):

did it work now?

OpenStudy (xlpoisonpotato):

hello?

OpenStudy (xlpoisonpotato):

anyone there?

OpenStudy (shadowlegendx):

Therefore angle CAD is 30 degrees

OpenStudy (xlpoisonpotato):

ok

OpenStudy (shadowlegendx):

ACD is 120 degrees

OpenStudy (xlpoisonpotato):

oh yeah

OpenStudy (xlpoisonpotato):

adc is acute

OpenStudy (xlpoisonpotato):

u still here

OpenStudy (shadowlegendx):

Read my message

OpenStudy (xlpoisonpotato):

ok

OpenStudy (xlpoisonpotato):

i still dont get it can u run me through the steps

OpenStudy (shadowlegendx):

It is explained better there

OpenStudy (shadowlegendx):

You just have to understand SOH CAH TOA

OpenStudy (shadowlegendx):

Sin Cos Tan

OpenStudy (xlpoisonpotato):

so you just go 400 over tan(30)

OpenStudy (xlpoisonpotato):

then the same over 60?

OpenStudy (xlpoisonpotato):

thats not working

OpenStudy (xlpoisonpotato):

OHH i get it'

OpenStudy (xlpoisonpotato):

you gotta take 400 divided by tan of 30 degrees minus 400 divided by the tan of 60 degrees to get 461.9

OpenStudy (xlpoisonpotato):

all who cheat off this is the future ur welcome i got it right

OpenStudy (mathstudent55):

You can solve this problem without trigonometry. Since the triangles are 30-60-90 right triangles, you can use the known ratios of the lengths of the sides. In a 30-60-90 triangle, the sides have lengths in the ratio: \(1 : \sqrt 3 : 2\) That is the ratio of the lengths of the short leg : long leg : hypotenuse In triangle ABD, AB is the short leg, and BD is the long leg, so \(BD = \sqrt 3 AC = 400 \sqrt 3 ~ft\approx 692.8 ~ft\) In triangle ABC, AB is the long leg, and BC is the short leg, so \(BC = \dfrac{400~ft}{\sqrt 3} \approx 230.9~ft\) Finally, \(CD = BD - BC = 692.8 ~ft- 230.9 ~ft= 461.9 ~ft\)